OpenStudy (anonymous):

solve each radical equation √5x+14=x

OpenStudy (anonymous):

take natural number to one side and square on both side

OpenStudy (anonymous):

what do you mean the natural number?

OpenStudy (anonymous):

\[sqrt{5x}-x=14 and square on each side\]

OpenStudy (anonymous):

14 = x - sqrt(5)x 14 = x(1-sqrt(5)) x = 14/(1-sqrt(5) x = around -11.33

OpenStudy (anonymous):

thanks!